# Constructing stable, high-order finite-difference operators on point clouds over complex geometries

### v1

@@ 5.4 Energy stability study
The final study is intended to verify the energy stability of the SBP discretization (24). We consider the annulus geometry $\Omega_{\text{ann}}$ with the spatially varying advection field
@@
We use $\beta=1/10$ to generate a quasi-uniform node distribution for the annulus, with $n_{r}=48$ and $n_{\theta}=288$ nodes in the radial and angular directions, respectively. We solve the ordinary differential equation (26) for one period, $T=2\pi$, using the classical fourth-order Runge Kutta method (RK4). Since RK4 is conditionally stable, we limit the time step to $\Delta t=2/\rho_{p}$, where $\rho_{p}$ is the spectral radius of the degree $p$ system matrix corresponding to (26).
@@ 6 Summary and future work
We conducted several numerical studies to verify the theoretical claims in the paper. Studies of the diagonal mass matrix demonstrated that the entries (i.e. quadrature weights) are well-behaved after solving the linear inequality, and that a solution exists for all but the coarsest node distributions. We also confirmed that the quadrature induced by the diagonal norm is order $2p$ for smooth functions. Finally, we used the SBP operators to discretize the linear advection equation and confirmed the accuracy and stability of the schemes.
There are numerous aspects of the point-cloud SBP operators that remain to be investigated, but we limit our discussion to just a few. Theoretically, we would like to see a more precise characterization of when the null space of $\mathsf{Z}^{T}$ is equal to the Vandermonde matrix $\mathsf{V}_{2p-1,\mathsf{d}}$. This could have implications for constructing node distributions and stencils that ensure the norm inequality is always feasible. Practically, we would like to use the operators on large-scale fluid-flow simulations, which will require some form of parallel implementation. Finally, we are interested in exploiting the flexibility of the point-cloud operators for solution adaptation and shape optimization.

### v2

@@ 5.5 Energy stability study
This study is intended to verify the energy stability of the SBP discretization (24). We consider the annulus geometry $\Omega_{\text{ann}}$ with the spatially varying advection field
@@
We use $\beta=1/10$ to generate a quasi-uniform node distribution for the annulus, with $n_{r}=48$ and $n_{\theta}=288$ nodes in the radial and angular directions, respectively. We solve the ordinary differential equation (27) for one period, $T=2\pi$, using the classical fourth-order Runge Kutta method (RK4). Since RK4 is conditionally stable, we limit the time step to $\Delta t=2/\rho_{p}$, where $\rho_{p}$ is the spectral radius of the degree $p$ system matrix corresponding to (27); $\rho_{p}$ is computed using the Julia interface to the ARPACK library [37].
@@
We conclude this section with Figure 12, which plots the absolute error distributions at the final time with and without dissipation. To conserve space, the figure shows the errors for $p=1$ and $p=4$ only; the results for $p=2$ and $p=3$ are qualitatively similar, but have error magnitudes between the $p=1$ and $p=4$ errors. The exact solution at node $\bm{x}_{i}$ and time $t=2\pi$ is approximated by numerically solving the ordinary differential equation $d\bm{x}/dt=\bm{\lambda}(\bm{x})$ with the terminal condition $\bm{x}(t=2\pi)=\bm{x}_{i}$, backward in time from $t=2\pi$ to $t=0$ using Verner’s 9/8 Runge-Kutta method [38] as implemented in the DifferentialEquations.jl Julia package [39], and then substituting the solution into $\mathcal{U}_{0}(\bm{x})$.
Figure 12 demonstrates the value of including some numerical dissipation with the point-cloud SBP discretizations; the solutions with dissipation contain fewer high-frequency oscillations compared to the solutions without dissipation. Furthermore, the $p=4$ errors are smaller in magnitude relative to the $p=1$ errors, as expected.
(a) $p=1$, no dissipation
(b) $p=1$, with dissipation
(c) $p=4$, no dissipation
(d) $p=4$, with dissipation
Figure 12: Absolute error distributions at the final simulation time $t=2\pi$ for $p=1$ (top) and $p=4$ (bottom) with no dissipation (left) and dissipation added (right).
5.6 Component runtimes and scaling
The paper’s final study investigates the relative computational cost and scaling of the components used to construct the point-cloud SBP operators. While our implementation has not been optimized, we hope the timings presented below provide some guidance regarding potential computational bottlenecks that will need to be addressed in future work.
We gather component runtimes using the airfoil geometry, $\Omega_{\text{foil}}$, and the linear advection problem described in Section 5.3. We consider the same distributions as the accuracy study, with point resolutions defined by $n_{y}\in\{4,8,16,32,64\}$ and $n_{x}=5n_{y}$. For each discretization, we time the following components in the construction process:
mesh:
Background mesh generation (Section 3.2).
stencil:
Stencil construction for each cell (Section 3.2.2).
$\mathsf{M}$ opt.:
Mass-matrix optimization (Section 4.3).
$\mathsf{S}_{x}$ $\mathsf{E}_{x}$:
Construction of $\mathsf{S}_{x}$ and $\mathsf{E}_{x}$ (Sections 3.3.2, 3.3.3, and 3.4).
$\mathsf{A}$ (diss.):
Construction of the dissipation operator (Section 3.6).
system:
Forming the discretization (Equation (24)).
While the list above is not exhaustive, the runtimes of these components were found to be representative of the overall costs.
Figure 13 shows the average component runtimes versus point resolution $n_{x}$, where the average times are based on ten distribution samples [^5] More precisely, we gather 11 samples and discard the first sample to avoid including Julia’s compile times.. The linear inequality fails for $p=4$ on the coarsest distribution ($n_{x}=20$) for all samples, so the average times for components that depend on the mass matrix are not shown for degree 4 and $n_{x}=20$. Note that a log scale is used for the axes, and a light gray reference is included in each subplot to show linear cost.
We highlight three observations based on Figure 13 and the component times more generally.
  • Using data from the two finest grids, we find that the components scale between $t\propto n_{x}^{1.2}$, for the background mesh construction, and $t\propto n_{x}^{1.5}$ for the $\mathsf{M}$ optimization problem. Thus, the times scale super linearly with $n_{x}$ and not linearly with the number of nodes ($N\propto n_{x}^{2}$), which is somewhat surprising. However, inspecting subcomponent timings reveals that the cost is driven by the number of cut cells, which scales with $n_{x}$ and not $N$. This is true even for the mass-matrix optimization, which is dominated by the time required to set up the linear inequality (20) and not the time to solve it.
  • Constructing $\mathsf{S}_{x}$ and $\mathsf{E}_{x}$ is the most expensive component at present. This can be attributed to the cut-cell related costs, particularly computing quadratures for the cut cells and their faces.
  • The system build — shown as brown diamonds in Figure 13 — is the fastest component for $p=1$ but the second slowest for $p=4$. This reflects the decrease in sparsity in the system matrix as $p$ increases, which then requires more operations when building the matrix.
Remark 9.
While we have performed limited three-dimensional simulations, our preliminary results suggest that the cut cells continue to dominate the component run times. Consequently, we expect the construction algorithm to scale at approximately $N^{2/3}$ in three dimensions.
Figure 13: CPU time, in seconds, for six representative components of the point-cloud SBP construction algorithm.
@@ 6 Summary and future work
We conducted several numerical studies to verify the theoretical claims in the paper. Studies of the diagonal mass matrix demonstrated that the entries (i.e. quadrature weights) are well-behaved after solving the linear inequality, and that a solution exists for all but the coarsest node distributions. We confirmed that the quadrature induced by the diagonal norm is order $2p$ for smooth functions. We used the SBP operators to discretize the linear advection equation and confirmed the accuracy and stability of the schemes. Our results also show that the point-cloud SBP operators are not as efficient as classical (tensor-product) finite-difference SBP operators. We investigated how the primary components of the construction algorithm scale and found that the cut-cells dictate the computational cost of most components.
There are numerous aspects of the point-cloud SBP operators that remain to be investigated, but we limit our discussion to just a few. Theoretically, we would like to see a more precise characterization of when the null space of $\mathsf{Z}^{T}$ is equal to the Vandermonde matrix $\mathsf{V}_{2p-1,\mathsf{d}}$. This could have implications for constructing node distributions and stencils that ensure the norm inequality is always feasible. Practically, we would like to use the operators on large-scale fluid-flow simulations, which will require some form of parallel implementation. To reduce computational cost, we are interested in exploiting the flexibility of the point-cloud operators for solution adaptation and shape optimization.
Finally, given that the present work builds on the generalized DGD method [18], future work could combine point-cloud SBP operators with radial-basis function (RBF) discretizations. For instance, RBFs could be incorporated into the accuracy conditions of the point-cloud SBP method. This has the potential to extend the energy-stable, global RBF-SBP method [40] by enabling a diagonal mass matrix and a sparse system matrix, i.e. an energy-stable, local RBF-FD method.
@@ Acknowledgements
Appendix A Theorem 4 Details
Here we provide additional steps showing how $\mathsf{Q}_{x}\mathsf{V}$ can be written as the five sums in Equation (13). First, we substitute the symmetric-skew-symmetric decomposition of $\mathsf{Q}_{x}$:
$\mathsf{Q}_{x}\mathsf{V}=\Big{(}\mathsf{S}_{x}+\frac{1}{2}\mathsf{E}_{x}\Big{)}\mathsf{V}$
Next, we substitute the global $\mathsf{S}_{x}$ and $\mathsf{E}_{x}$ definitions given by Equations (12) and (11), respectively.
$\mathsf{Q}_{x}\mathsf{V}=\sum_{\mathsf{c}=1}^{N_{C}}\big{(}\mathsf{P}^{\mathsf{c}}\big{)}^{T}\;\mathsf{S}_{x}^{\mathsf{c}}\;\mathsf{P}^{\mathsf{c}}\;\mathsf{V}+\frac{1}{2}\sum_{\mathsf{f}\in F_{I}}\big{(}\mathsf{R}^{\mathsf{f},-}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{R}^{\mathsf{f},+}\;\mathsf{V}\\ -\frac{1}{2}\sum_{\mathsf{f}\in F_{I}}\big{(}\mathsf{R}^{\mathsf{f},+}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{R}^{\mathsf{f},-}\;\mathsf{V}+\frac{1}{2}\sum_{\mathsf{f}\in F_{B}}\big{(}\mathsf{R}^{\mathsf{f}}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{R}^{\mathsf{f}}\;\mathsf{V}$ (29)
In the first sum above, the skew-symmetric operator for cell $\mathsf{c}$ can be expressed as $\mathsf{S}_{x}^{\mathsf{c}}=\mathsf{Q}_{x}^{\mathsf{c}}-\frac{1}{2}\mathsf{E}_{x}^{\mathsf{c}}$. Using this identity, and the interpolated Vandermonde matrices defined in the Proof of Theorem 4, Equation (29) becomes
$\mathsf{Q}_{x}\mathsf{V}=\sum_{\mathsf{c}=1}^{N_{C}}\big{(}\mathsf{P}^{\mathsf{c}}\big{)}^{T}\;\mathsf{Q}_{x}^{\mathsf{c}}\;\mathsf{V}^{\mathsf{c}}-\frac{1}{2}\sum_{\mathsf{c}=1}^{N_{C}}\big{(}\mathsf{P}^{\mathsf{c}}\big{)}^{T}\;\mathsf{E}_{x}^{\mathsf{c}}\;\mathsf{V}^{\mathsf{c}}+\frac{1}{2}\sum_{\mathsf{f}\in F_{I}}\big{(}\mathsf{R}^{\mathsf{f},-}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{V}^{\mathsf{f}}\\ -\frac{1}{2}\sum_{\mathsf{f}\in F_{I}}\big{(}\mathsf{R}^{\mathsf{f},+}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{V}^{\mathsf{f}}+\frac{1}{2}\sum_{\mathsf{f}\in F_{B}}\big{(}\mathsf{R}^{\mathsf{f}}\big{)}^{T}\;\mathsf{B}^{\mathsf{f}}\;\mathsf{N}_{x}^{\mathsf{f}}\;\mathsf{V}^{\mathsf{f}}.$
The five sums on the right-hand side above are equal to the five sums on the right-hand side of (13).